OpenStudy (anonymous):

what is this equation? 3(x-7)-x=2x-21 7 -2 infinite number of solutions no solution

OpenStudy (carolina4567):

3x%2B7-x=2x%2B21 Start with the given equation. 2x%2B7=2x%2B21 Combine like terms on the left side. 2x=2x%2B21-7 Subtract 7 from both sides. 2x-2x=21-7 Subtract 2x from both sides. 0x=21-7 Combine like terms on the left side. 0x=14 Combine like terms on the right side. 0=14 Simplify. Since this equation is NEVER true for any x value. So there are no solutions.

OpenStudy (carolina4567):

Dont mind the percentage its a glitch

OpenStudy (carolina4567):

Therefore D :)

OpenStudy (anonymous):

i think my brain stopped workin for a moment XD